OpenStudy (anonymous):

y+5=7/4x Identify point (x.y) and slope y+6=(x-1) Identify point (x.y) and slope

OpenStudy (anonymous):

y+5=7/4x subtract the 5 from both sides You are left with Y=7/4x-5 Using the slope intercept formula y=mx+b we can determine that the slope is 7/4 m=7/4

OpenStudy (anonymous):

Which one is x and y?

OpenStudy (anonymous):

Slope for x is 1.750 Slope for y is 1.000